# LightEndoStereo: A Real-time Lightweight Stereo Matching Method for Endoscopy Images

### v1

@@ III-F Loss Function
$\begin{split}\text{Loss}=w_{1}\cdot\|d_{f}-d_{gt}\|_{1}+w_{2}\cdot\|d_{cg}-d_{gt}\|_{1}+\\ w_{3}\cdot\|d_{dr}-d_{gt}\|_{1},\end{split}$ (14)
@@
Figure 3: Visualization of disparity estimation on SCARED and SERV-CT datasets. (a) and (b) are from SCARED, while (c) and (d) are from SERV-CT. When overall depth variation is minimal, as in (b), most models perform similarly. However, in scenes with significant depth transitions—indicating high-frequency components, such as between surgical instruments and tissues in (a)—LightEndoStereo provides more accurate depth estimates at boundary regions. Figures (c) and (d) show that LightEndoStereo also performs well on SERV-CT.
IV Experiments
IV-A Datasets
We trained our algorithm on the SCARED dataset and tested it on both the SCARED and SERV-CT datasets.
IV-A1 SCARED Dataset [24]
@@ IV-A2 SERV-CT Dataset [26]
IV-B Implementation details
We adopted a MobileNetV4, pre-trained on ImageNet, as the decoder part branch of feature network. The model was implemented using the PyTorch framework. For optimization, we utilized the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate for the feature network is $1\times 10^{-4}$ and $1\times 10^{-3}$ for the other parts. During training, we set the random seed to 6. For data augmentation, we applied random cropping to resize the training images to $256\times 512$. The model was trained for 100 epochs on the SCARED dataset. For evaluation, we tested the model on datasets 8 and 9 of SCARED, as well as the SERV-CT dataset. All experiments were conducted on an Ubuntu 22.04 system with 4 Nvidia RTX 2080Ti GPUs.
IV-C Evaluations on SCARED and SERV-CT
TABLE I: MAE (unit: mm) for the SCARED Test Set. Each test set comprises 5 keyframes (kf1–kf5). Lower values indicate better performance. Methods marked with (*) reference results from other papers. Bold font indicates the optimal results.
@@ IV-C Evaluations on SCARED and SERV-CT
GwcNet [16] 9.076 2.926 1.433 1.7 1.203 3.868 1.138 2.934 1.997 0.6135 2.689
PSMNet* [29] 8.96 2.77 1.43 1.83 1.42 3.99 1.08 2.82 1.95 0.56 2.68
RAFT [21] 7.921 2.362 1.721 2.167 1.88 4.363 1.087 2.872 1.983 1.549 2.79
IGEV [20] 7.955 2.393 1.697 2.265 1.938 3.893 1.174 2.895 2.504 1.25 2.796
DLNR [19] 7.848 2.529 2.654 3.209 3.418 5.374 1.913 4.324 4.82 10.25 4.634
Selective-Stereo [31] 7.537 2.409 1.668 2.31 2.264 3.86 0.958 2.838 1.775 1.018 2.664
MSDESIS-resnet32 [7] 10.14 3.467 3.306 3.889 2.099 5.428 1.636 3.378 1.768 0.6375 3.575
MSDESIS-light [7] 11.45 15.91 7.727 10.67 9.849 98.39 130.8 66.57 101.8 203.4 65.67
DCStereo* [34] 8.38 2.77 1.54 1.82 1.19 3.99 1.09 2.64 2.09 1.04 2.65
LightEndoStereo(Ours) 7.62 2.316 1.669 2.281 1.025 4.428 1.167 2.809 2.125 0.4793 2.592
TABLE II: Results of comparison experiments on the SERV-CT dataset.Methods marked with (*) reference results from other papers. Bold font indicates the optimal results.
Model MAE(pixel)$\downarrow$ D1$\downarrow$ Bad1$\downarrow$ Bad2$\downarrow$ Bad3$\downarrow$
GwcNet [16] 5.813 33.92% 79.37% 56.32% 40.93%
PSMNet* [29] 6.355 - 70.88% 49.80% 33.87%
RAFT [21] 21.500 87.50% 95.80% 91.80% 88.50%
IGEV [20] 32.895 92.91% 98.28% 96.53% 94.76%
DLNR [19] 22.030 88.97% 96.21% 92.61% 89.13%
Selective-Stereo [31] 24.103 86.36% 95.68% 91.33% 87.01%
MSDESIS-resnet32 [7] 5.742 28.26% 78.81% 55.44% 37.79%
MSDESIS-light [7] 7.065 41.35% 81.09% 63.17% 49.22%
LightEndoStereo(Ours) 2.367 14.34% 58.59% 31.52% 18.49%
We evaluated our algorithm against several methods for natural image stereo matching and two algorithms for endoscopic data. Specifically, GwcNet [16], PSMNet [29], RAFT [35], IGEV [20], DLNR [19], and Selective-Stereo [31] are methods designed for natural images, with DLNR and Selective-Stereo focusing on optimizing high-frequency details in disparity maps. For these natural image methods, except for PSMNet, we retrained them on the SCARED dataset. For PSMNet, we directly cited the results from [34, 3]. MSDESIS [7] and DCStereo [34] are algorithms tailored for endoscopic images. DCStereo’s evaluation methodology on SCARED aligns with ours, so we directly cited their results. However, since DCStereo was not tested on the SERV-CT dataset and its code is not publicly available, we omitted their results on SERV-CT. MSDESIS was trained on the SCARED dataset, and we directly loaded their pre-trained weights. Given that our maximum disparity is set to 192, whereas MSDESIS uses a maximum disparity of 320, the results we reproduced differ slightly from those reported in their paper. The evaluation results on SCARED are shown in Table I, and the results on SERV-CT are shown in Table II. LightEndoStereo achieved SOTA performance on both datasets and demonstrated excellent generalization on SERV-CT.
IV-D Runtime Evaluation
TABLE III: Quantitative results of model computational performance. Note that only methods with an average MAE < 3mm on the SCARED dataset are included.Methods marked with (*) reference results from other papers. Bold font indicates the optimal results.
@@ IV-D Runtime Evaluation
GwcNet [16] 6.43 2.466 26.4
PSMNet* [29] 3.672 - 225
RAFT [21] 11.11 3.757 468.6
IGEV [20] 12.5 3.288 434.5
Selective-Stereo [31] 13.141 4.062 516.52
DCStereo* [34] 3.404 - 191
LightEndoStereo(Ours) 11.094 0.846 23.38
To evaluate the runtime performance of our model, we conducted tests using a single Nvidia RTX 2080Ti GPU, paired with an Intel(R) Xeon(R) Gold 6130T CPU @ 2.10GHz. We sampled 100 images ($1280\times 1024$) from dataset 8 of the SCARED dataset for testing and used the average inference time as the evaluation metric. The results are shown in Table III. Although the MSDESIS-Light model has the smallest footprint, its matching accuracy is poor. In contrast, LightEndoStereo achieved a real-time matching speed of 23.38 ms per frame.
@@
TABLE IV: Ablation Study Results. We use the average MAE on the SCARED dataset as the accuracy evaluation metric. When ablating the 3D UNet-MCA, we employ the widely-used 3D Hourglass aggregation network.
3D UNet-MCA HFDO MAE(mm) Runtime(ms) Params(M) FLOPs(T)
2.718 25.412ms 13.0479 1.7552
$\checkmark$ 2.645 25.972ms 13.0773 1.7579
2.621 22.655ms 11.0649 0.8435
$\checkmark$ 2.592 23.386ms 11.0943 0.8461
IV-E Ablation Study
We conducted detailed ablation studies on the proposed 3D MCA and HFDO modules (see Table IV). When ablating the 3D UNet-MCA aggregation network, we employ the widely-used 3D Hourglass aggregation network [16]. The 3D Hourglass aggregation network is a stacked symmetric multi-scale network that features a large number of parameters and high computational complexity. As shown in Table III, the 3D UNet-MCA not only reduces the number of parameters and computational complexity but also enhances disparity estimation accuracy. As Figure 4 illustrates, the HFDO module significantly reduces matching errors at tissue boundaries, leading to notable improvements in both the Stacked 3D Hourglass and 3D UNet-MCA configurations.
V Conclusions and limitations
We introduce LightEndoStereo, a novel lightweight real-time stereo matching algorithm tailored for endoscopic images. Our approach integrates a coordinate attention mechanism and the Mamba block to achieve efficient cost aggregation via a lightweight 3D UNet. Moreover, by enhancing high-frequency features in the wavelet domain, we optimize disparity estimation, effectively mitigating large matching errors at tissue boundaries—a common challenge in endoscopic stereo matching. During validation, LightEndoStereo exhibited strong generalization on the SERV-CT dataset, a potential that warrants further exploration. Acknowledging the significance of generalizable stereo matching algorithms, we intend to delve deeper into this area in our future research.

### v3

@@ 3.5 Loss Function
$\begin{split}\text{Loss}=w_{1}\cdot\|d_{f}-d_{gt}\|_{1}+w_{2}\cdot\|d_{cg}-d_{gt}\|_{1}+w_{3}\cdot\|d_{dr}-d_{gt}\|_{1},\end{split}$ (12)
@@
4 Experiments
4.1 Datasets
We evaluate our method on two publicly available dataset:
@@ 4.1 Datasets
4.2 Implementation Details
Our implementation is based on the PyTorch framework. We utilized the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is set to $1\times 10^{-4}$. For data augmentation, we applied random cropping to resize the training images to $256\times 512$. The model was trained for 100 epochs on the SCARED dataset. For evaluation, we tested the model on datasets 8 and 9 of the SCARED dataset, as well as the entire SERV-CT dataset. All experiments were conducted on an Ubuntu 22.04 system with 4 Nvidia RTX 2080Ti GPUs.
4.3 Evaluations on SCARED and SERV-CT
We evaluated our algorithm against a set of state-of-the-art stereo matching methods, including both those designed for natural images and those specifically developed for endoscopic data. In particular, GwcNet [10], PSMNet [3], RAFT [17], IGEV [28], DLNR [35], and Selective-Stereo [26] are representative natural image methods. Among them, DLNR and Selective-Stereo explicitly focus on optimizing high-frequency details in disparity estimation. Except for PSMNet, which we report using results cited from [13, 27], all other models were retrained on the SCARED dataset under a consistent training pipeline.
Table 1: Results of comparison experiments on the SERV-CT dataset.
[…]
GwcNet 5.813 33.92% 79.37% 56.32% 40.93%
PSMNet* 6.355 - 70.88% 49.80% 33.87%
RAFT 21.500 87.50% 95.80% 91.80% 88.50%
IGEV 32.895 92.91% 98.28% 96.53% 94.76%
DLNR 22.030 88.97% 96.21% 92.61% 89.13%
Selective-Stereo 24.103 86.36% 95.68% 91.33% 87.01%
MSDESIS-full 5.742 28.26% 78.81% 55.44% 37.79%
MSDESIS-light 7.065 41.35% 81.09% 63.17% 49.22%
RRESM(ours) 2.367 14.34% 58.59% 31.52% 18.49%
Table 2: MAE (unit: $m$) for the SCARED Test Set. Each test set comprises 5 keyframes (kf1–kf5). Lower values indicate better performance. Methods marked with (*) reference results from other papers. Bold indicates the best performance, and underlined values indicate the second-best.
@@ 4.3 Evaluations on SCARED and SERV-CT
GwcNet 9.07 2.92 1.43 1.70 1.20 3.86 1.13 2.93 1.99 0.61 2.68
PSMNet* 8.96 2.77 1.43 1.83 1.42 3.99 1.08 2.82 1.95 0.56 2.68
RAFT 7.92 2.36 1.72 2.16 1.88 4.36 1.08 2.87 1.98 1.54 2.79
IGEV 7.95 2.39 1.69 2.26 1.93 3.89 1.17 2.89 2.50 1.25 2.79
DLNR 7.84 2.52 2.65 3.20 3.41 5.37 1.91 4.32 4.82 10.25 4.63
Selective-Stereo 7.53 2.40 1.66 2.31 2.26 3.86 0.95 2.83 1.77 1.01 2.66
MSDESIS-full 10.14 3.46 3.30 3.88 2.09 5.42 1.63 3.37 1.76 0.63 3.57
MSDESIS-light 11.45 15.91 7.72 10.67 9.84 98.39 130.80 66.57 101.80 203.40 65.67
DCStereo* 8.38 2.77 1.54 1.82 1.19 3.99 1.09 2.64 2.09 1.04 2.65
RRESM(ours) 7.62 2.31 1.66 2.28 1.02 4.42 1.16 2.80 2.12 0.47 2.59
In contrast, MSDESIS [20] and DCStereo [13] are methods tailored for endoscopic stereo matching. For DCStereo, we directly cited its results on the SCARED dataset, as their evaluation protocol is consistent with ours. However, since DCStereo was not evaluated on the SERV-CT dataset and its implementation is not publicly available, we omitted it from SERV-CT comparisons. As for MSDESIS, we used the pre-trained model provided by the authors. Note that our method uses a maximum disparity of 192, while MSDESIS adopts 320, which may account for slight differences between our reproduced results and those reported in their original paper.
Figure 5: Visualization of disparity estimation on the SCARED and SERV-CT datasets. (a) and (b) are from SCARED, while (c) and (d) are from SERV-CT. In cases with limited depth variation, such as (b), most methods perform similarly. However, in high-frequency regions like surgical tool-tissue boundaries in (a), RRESM yields more accurate depth predictions. On SERV-CT, our method also delivers competitive results. Note: MAE is measured in millimeters ($m$) for SCARED (with ground-truth depth) and in pixels for SERV-CT (with ground-truth disparity).
The quantitative results on the SCARED and SERV-CT datasets are presented in Table 2 and Table 1, respectively. Sample visualizations are shown in Fig. 5. On the SCARED dataset, RRESM achieves the best average MAE of 2.59 mm, outperforming both natural image and endoscopy-specific baselines. On the SERV-CT dataset, our method also achieves superior performance across all metrics, including the lowest MAE (2.367 pixels), lowest Bad-n error rates, and a significantly lower D1 score. These results highlight RRESM’s robustness in handling challenging tissue boundaries and its strong generalization ability to unseen clinical data.
4.4 Runtime Evaluation
Table 3: Quantitative results of model computational performance.
@@ 4.4 Runtime Evaluation
GwcNet 6.43 2.466 26.4
PSMNet* 3.672 - 225
RAFT 11.11 3.757 468.6
IGEV 12.5 3.288 434.5
Selective-Stereo 13.141 4.062 516.52
DCStereo* 3.404 - 191
RRESM(ours) 11.094 0.846 23.38
To evaluate the runtime performance of our model, we conducted tests using a single Nvidia RTX 2080Ti GPU. We sampled 100 images ($1280\times 1024$) from dataset 8 of the SCARED for testing and used the average inference time as the evaluation metric. The results are shown in Table 3. Although the MSDESIS-Light model has the smallest footprint, its matching accuracy is poor. In contrast, RRESM achieved a real-time matching speed of 23.38 ms per frame.
4.5 Ablation Study
Table 4: Ablation Study Results. The average MAE on the SCARED dataset is used as the accuracy evaluation metric.
MCA HFDO MAE($m$) Runtime(ms) Params(M) FLOPs(T)
2.718 25.412ms 13.047 1.755
$\checkmark$ 2.645 25.972ms 13.077 1.757
2.621 22.655ms 11.064 0.843
$\checkmark$ 2.592 23.386ms 11.094 0.846
We conducted ablation studies to evaluate the contributions of the proposed 3D Mamba Coordinate Attention (MCA) and High-Frequency Disparity Optimization (HFDO) modules. MCA is designed to improve cost aggregation efficiency by modeling axis-specific long-range dependencies, while HFDO focuses on refining depth estimates in regions with high-frequency disparity transitions, such as tissue-tool boundaries.
@@ 4.5 Ablation Study
The experiments were performed on the SCARED dataset using our 3D U-Net-based architecture. To assess MCA, we replaced it with a conventional stacked 3D Hourglass network [10], which is commonly used in existing stereo matching pipelines. For evaluating HFDO, we simply removed the wavelet-based refinement module from the full pipeline. We report the Mean Absolute Error (MAE) in millimeters, runtime per frame, model parameters, and FLOPs, as summarized in Table 4.
Results show that both modules individually contribute to improved accuracy. Incorporating MCA alone reduces MAE from 2.718 to 2.621 mm while significantly lowering model complexity. Adding HFDO further improves performance to 2.592 mm. Notably, the complete model achieves the best trade-off between accuracy and efficiency. As visualized in Fig. 6, HFDO particularly enhances boundary precision by reducing local matching errors in high-frequency regions. These results validate the effectiveness and complementary nature of MCA and HFDO within the overall architecture.
5 Discussion and Conclusions
We present RRESM, a real-time stereo matching framework tailored for endoscopic imagery. By integrating a 3D Coordinate Attention mechanism with the Mamba block, RRESM enables efficient and lightweight cost aggregation while capturing long-range dependencies. The High-Frequency Disparity Optimization module further enhances disparity accuracy near anatomical edges through wavelet-based refinement. Extensive experiments on the SCARED and SERV-CT datasets show that RRESM achieves state-of-the-art performance and generalizes well to unseen domains, running at 42 FPS on high-resolution inputs.
Despite these promising results, our method has several limitations. Its performance may degrade under extreme illumination variation, specular reflections, or endoscope lens occlusion. Additionally, the current pipeline is limited to stereo input and assumes spatial rectification. In future work, we will (1) explore uncertainty-aware disparity refinement, (2) extend RRESM to multi-view and monocular depth estimation, and (3) validate its robustness across diverse anatomical sites and surgical conditions.